OpenStudy (anonymous):

The speed of light is 186,000 miles per second. What is this speed in scientific notation?

OpenStudy (xxdragoonxx):

Divide the 186,000 by 10, 2 time(s), making it a 1.86. 186,000 / 10 = 18600 / 10 = 1.86 Each division by 10 moves the decimal point to the left one digit. But in order to keep our number the same number, we must have a × 102 next to the 1.86. So, the final answer is \[1.86 × 102^2\]

OpenStudy (anonymous):

thanks

OpenStudy (xxdragoonxx):

Your welcome.

OpenStudy (anonymous):

How do you write the following expanded number in standard notation? 500 + 80 + 3 + 0.5 + 0.02